OpenStudy (anonymous):

Write 32x^5y^5 with only one exponent. Use parentheses.

OpenStudy (anonymous):

32=2^5

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

What about the parentheses?

OpenStudy (anonymous):

32x^5y^5 = 2^5 x^5 y^5 = (2xy)^5

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yep so i just write the very last part as my answer

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Thank you